OpenStudy (anonymous):

The gum you like to buy is on sale. It is regularly priced at $1.59 write an equation that will help you determine how much you'll ave if you buy the pack today s=sale price c=cost savings

OpenStudy (anonymous):

1.59 - s = c

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no problem...glad to help :)

OpenStudy (anonymous):

can you help me with a few more?

OpenStudy (anonymous):

possibly...what you got ?

OpenStudy (anonymous):

wx/y = z I am not sure about this...let me get someone here who might know...hold on

OpenStudy (anonymous):

@Mertsj ...can you help

OpenStudy (anonymous):

@hartnn am I doing this right ?

OpenStudy (mertsj):

c=1.59-s

OpenStudy (mertsj):

If you want to know how much you will save, take the regular price and subtract the sale price. For example: If a CD normally costs $10 and today it is on sale for $8 you find your savings by doing: 10-8

OpenStudy (anonymous):

what about the second problem ? Did I do that one correctly ?

OpenStudy (mertsj):

yes. wx/y=z is correct.

OpenStudy (anonymous):

oh thank you so much :)

OpenStudy (anonymous):

thanks can you help me with 2 more?

OpenStudy (anonymous):

I can try :)

OpenStudy (anonymous):

jaun works at a retail store he earn $6 an hour and earns a $5 bonus everytime he get someone to sign up for a in sstore credit card write a formula to help juan determine how much he will earn in a week a=total amount he earned h=hours worked in one week b=number of cards sold in one week

OpenStudy (anonymous):

6h + 5b = a

OpenStudy (anonymous):

okay that was right just one more

OpenStudy (anonymous):

james just bought a large container of ice cream the label says that there are 16 serving of ice cream write a formula to help james determine the number of gram sin one serving of ice cream x=number of grams of ice cream in container y=number of grams per seving

OpenStudy (anonymous):

y x 16 = x

OpenStudy (anonymous):

or 16y = x

OpenStudy (anonymous):

hartnn...did I do this one correctly ?

OpenStudy (anonymous):

I think I did blah105

OpenStudy (anonymous):

which one?

hartnn (hartnn):

yes, correct

OpenStudy (anonymous):

the last one....I think it can be written either way

OpenStudy (anonymous):

thanks hartnn...I just needed to check

OpenStudy (anonymous):

blah...hartnn is very good, and he can help many people :)

OpenStudy (anonymous):

It was wrong can you help me with the last one i didnt know there was 5

OpenStudy (anonymous):

it was wrong ?

OpenStudy (anonymous):

Thats what it said but it okay

hartnn (hartnn):

i think the equation is correct, but we needed y so , y= x/16

OpenStudy (anonymous):

oh...I see

OpenStudy (anonymous):

kayla notices that she can buy snacks and candy in bulk at the grocery sstore the snacks she wants to buy weigh 1 kilogram write a formula that will help kayla determine how much shell have to pay if she buys food in in bulk t=total price p=price per kilogram

OpenStudy (anonymous):

p = t ??

OpenStudy (anonymous):

She is only buying one kilogram, so the price per kilogram is the price she will pay ? right ?

OpenStudy (anonymous):

it was right thanks so much you've been a big help:)

OpenStudy (anonymous):

your welcome...I tried my best